12. You have $10,000 in savings. Assume you earn 2.5% compounded quarterly on it for three years, then you earn 5% annually for the next two years, then you withdraw $1,000 to buy a new phone, then you earn 7% per year on the remaining balance for 10 more years. g. How much savings will you have at the end of the entire time? h. How much would you have had if you had not purchased the phone?
Answer #1

Present value = Future value/(1+i)^n

i = interest rate per period

n= number of periods

g)

Future value before withdraw

= 10000 * (1+ 0.025/4)^12 * (1+0.05)^2

= 11880.90

total savings

= (11880.90 - 10000) * (1+7%)^10

= 3700.02

h)

total savings =

10000 * (1+ 0.025/4)^12 * (1+0.05)^2 * (1+7%)^10

= 23371.53